OpenStudy (anonymous):

Solve: 2(x + 8) − (x + 1) = 20

OpenStudy (jlastino):

2(x + 8) − (x + 1) = 20 2x+16-x-1=20 x+15=20 x=5

OpenStudy (anonymous):

2x+16-1x+-1=20

OpenStudy (anonymous):

X+15=20

OpenStudy (anonymous):

X=5